OpenStudy (anonymous):

The circle above has a radius of 11 cm. What is the area of the circle? Use = 3.14

OpenStudy (anonymous):

the area of a circle is \[\pi r^2\] we know that r = 11 we know that pi = 3.14 so the answer is 3.14 * 11^2 = 3.14 * 121

OpenStudy (anonymous):

The circumference of the bike tire above is 83.524 inches. What is the radius of the bike tire? (Use 3.14 for .)

OpenStudy (anonymous):

circumference = 2*pi*r pi = 3.14 r = circumference/(2pi) r = 83.524/(2*3.14)

OpenStudy (anonymous):

volume = width * length * height V = 3*3*3 +3*3*6

OpenStudy (anonymous):

ok 2 more thanks